The following financial information is for Alpha Corporation are for the fiscal years ending 2017 & 2016 (all balances are normal):

Item/Account 2017 2016
Accounts Receivable $9800 $9400
Inventory $30000 $16000
Net Sales (all credit) $85000 $71000
Cost of Goods Sold $62000 $52000
Net Income $7200 $6900

Use this information to determine for FY 2017: (Round & enter your answers to one decimal place and enter the value.)

1. the inventory turnover ratio

2. number of days of inventory

3. Gross Profit Margin

4. Profit Margin
Answer #1

Inventory Turnover ratio = Cost of Goods Sold / Average inventory

= 62000 / [(30,000+16,000)/2]

= 2.70 (rounded)

Days in inventory = 365/ Inventory Turnover

= 365/2.70

= 135.2 days

Gross profit margin = (Sales - Cost of Goods Sold) / sales

= (85,000-62,000)/85,000

= 27.1%

Profit Margin = Net income / sales

= 7200/85000

= 8.5%

I have rounded some numbers.

Please comment if you face any issues